OpenStudy (diana):

How many years will it take for an initial investment of $10,000 to grow to $25,000? Assume a rate of interest of 6% compounded continuously.

OpenStudy (anonymous):

25000=10000e^(0.06t)

OpenStudy (diana):

but how do you solve for t?

OpenStudy (anonymous):

you can do it y using log on both sides

OpenStudy (diana):

so \[t = \ln (25000)\div \left[ \ln (10000) \times(.06) \right]\]?

OpenStudy (anonymous):

ys